In the book Advanced Managerial Accounting, Robert P. Magee discusses monitoring cost variances. A cost variance is the diffez -3.9 -3.8 -3.7 -3.6 -3.5 -3.4 -3.3 -3.2 -3.1 -3.0 -2.9 -2.8 -2.7 -2.6 -2.5 -2.4 -2.3 -2.2 -2.1 -2.0 -1.9 -1.8 -1.7 -1.6 -1.

Please help me solve this question using the table listed underneath the question.
Answer #1

(a)

µ 5 µ 1
σ 5353 σ 10112
x 2911 x 2911
z = (x - µ)/σ 0.54287316 z = (x - µ)/σ 0.2877769
P(z > 0.54) 0.2936 P(z > 0.29) 0.3868
Process A 0.2936
Process B 0.3868


Process B  is investigated more often.

(b)

µ 7641 µ 6721
σ 5353 σ 10112
x 2911 x 2911
z = (x - µ)/σ -0.88361666 z = (x - µ)/σ -0.37678006
P(z > -0.88) 0.8115 P(z > -0.38) 0.6468
Process A 0.8115
Process B 0.6468

Process A is investigated more often


(c) Process B is investigated more often.

(d)

µ 6721
σ 10112
µ 1 x 5461.436487
σ 10112
z 0.539995697 z = (x - µ)/σ -0.12456127
k = z*σ + µ 5461.436487 P(z > -0.16) 0.5496


Please give me a thumbs-up if this helps you out. Thank you!